if inflation is 2 percent a year compounded annually what would it cost in 13 years to buy a house currently valued at 75000.00.
Answer by lwsshak3(11628)

if inflation is 2 percent a year compounded annually what would it cost in 13 years to buy a house currently valued at 75000.00.
**
The formula to use: A=P(1+i)^n, A=amount after n periods, P=initial investment, i=interest per period, n=number of periods. This is callled the compound interest formula.
For given problem: A=to be determined, P=75000, i=.02, n=13
A=75000(1+.02)^13
A=75000(1.02)^13
A=97020
ans:
In 13 years it would cost $97020.50 to buy a house currently valued at $75000.00.